Financial data for Joel de Paris, Inc., for last year follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 126,000 $ 130,000
Accounts receivable 348,000 473,000
Inventory 567,000 482,000
Plant and equipment, net 819,000 815,000
Investment in Buisson, S.A. 402,000 430,000
Land (undeveloped) 250,000 251,000
Total assets $ 2,512,000 $ 2,581,000
Liabilities and Stockholders' Equity
Accounts payable $ 385,000 $ 337,000
Long-term debt 970,000 970,000
Stockholders' equity 1,157,000 1,274,000
Total liabilities and stockholders' equity $ 2,512,000 $ 2,581,000


Joel de Paris, Inc.
Income Statement
Sales $ 3,948,000
Operating expenses 3,237,360
Net operating income 710,640
Interest and taxes:
Interest expense $ 121,000
Tax expense 201,000 322,000
Net income $ 388,640


The company paid dividends of $271,640 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Do not round intermediate calculations and round your final answers to 2 decimal places.)

3. What was the company’s residual income last year?
Answer #1

PagenoO Anşlen- J, The companyâ ave siage opematimg Asgets. Begimning opeoiat ing a §ãet§ = 2,512,000 - 402,00 -250,000 1860Pagemo@ 2, The Componyş margim twmoven, and retunn on imveştmunt CROI) Net operatingncome Margin వయి F10,640 3৭५৪, ০०0 HanmnPageNo@ 3 he Reşidual imcome Company Reşidualîncome 154.) = t10, 640 -(1880000x15% = 410,640 - 282000 428640 ; Reşidual încom